OpenStudy (anonymous):

Carbon-14 has a half-life of 5730 years. In a plant fossil, you find that the 14C has decayed to 1/8.00 of the original amount. How long ago was this plant alive?

OpenStudy (anonymous):

hello. eq: Nf = Ni*(1/2)^t/t1/2 Nf = 1/8 Ni = 1 we have to find t so : 1/8 = 1*2^(-t/5730years) 2^-3 = 2 ^ (-t/5730) from there: -3 = -t/5730=>t = 5730*3=17190years

OpenStudy (anonymous):

Thank u